Saturday, April 2, 2011

Day 4, 04/02/11 -- Epoxy, for example

In case you hadn't realized, yesterday's post was a (lame) April Fool's joke. The Helium-10 fridge is fine, and everything is moving along.

Yesterday's update was basically more progress on all fronts. Kyle continued to put the instrument together, I finished the housekeeping wiring feedthrough wire replacement and made a new charcoal getter for the cryostat while Jeff finished soldering his wiring. Like much of our wiring, Jeff potted the connectors, which means he encapsulated the wires at the back of the connectors with epoxy (in this case, Stycast 2850FT, an alumina-filled epoxy designed for relatively high thermal conductivity and a metal-like thermal expansion coefficient) in order to prevent wires from moving and shorting together. We typically use the 2850FT resin with a low-viscosity catalyst, which is great for getting the epoxy to flow between wires, around contacts, and into the nooks and crannies (sort of like an english muffin) of the mold. Unfortunately for Jeff, the connectors he used on one end of his wiring, for some reason, had somewhat loose contacts, and the epoxy flowed around them and into the front face of the connector, rendering it unusable.

Today we decided to cut off the bad connectors and replace them. Luckily, unlike most of the connectors in the experiment, these particular connectors (25-pin D-subminiature connectors, the ubiquitous DB-25) are common and we were able to get replacements from the local Radio Shack. These connectors seem to have much more firmly-embedded contacts, so we don't expect the same problem to occur.

On other fronts, Kyle continued with the instrument -- he finished up making all of the sub-Kelvin thermal connections and we started on the process of optics alignment. This is a tricky task with tight tolerances: 0.005" in translation and 0.1 degree in angle. In order to measure these distances, we use a neat tool called a Microscribe MX, which is an articulated arm used to measure locations of...basically anything. That process continues.

In the meantime, I worked on some internal cryostat wiring while Jeff removed the components in the so-called "optics stack" of the instrument, which holds various optical filters and components, in preparation for upcoming alignment tasks.

Unfortunately, I forgot my camera's memory card and card reader in the high bay, so I have no pictures from today (I only took a few anyway). You'll have to make do with the pictures from yesterday that I didn't upload. I'll edit tomorrow with the link to today's pictures.

Yesterday's pictures:
https://picasaweb.google.com/asad137/EBEXInPalestine20110401

EDIT: Today's pictures:
https://picasaweb.google.com/asad137/EBEXInPalestine20110402

No comments:

Post a Comment